# Train with Perturbation, Infer after Merging: A Two-Stage Framework for Continual Learning

### v3

@@ Abstract
Continual Learning (CL) aims to enable models to continuously acquire new knowledge from a sequence of tasks with avoiding the forgetting of learned information. However, existing CL methods only rely on the parameters of the most recent task for inference, which makes them susceptible to catastrophic forgetting. Inspired by the recent success of model merging techniques, we propose Perturb-and-Merge (P&M), a novel continual learning framework that integrates model merging into the CL paradigm to mitigate forgetting. Specifically, after training on each task, P&M constructs a new model by forming a convex combination of the previous model and the newly trained task-specific model. Through theoretical analysis, we minimize the total loss increase across all tasks and derive an analytical solution for the optimal merging coefficient. To further improve the performance of the merged model, we observe that the degradation introduced during merging can be alleviated by a regularization term composed of the task vector and the Hessian matrix of the loss function. Interestingly, we show that this term can be efficiently approximated using second-order symmetric finite differences, and a stochastic perturbation strategy along the task vector direction is accordingly devised which incurs no additional forward or backward passes while providing an effective approximation of the regularization term. Finally, we combine P&M with LoRA, a parameter-efficient fine-tuning method, to reduce memory overhead. Our proposed approach achieves state-of-the-art performance on several continual learning benchmark datasets.
@@ 3.1 Infer after Merging
where $\hat{p}_{y}$ is the predicted probability for the correct class $y$ and $\mathbb{I}(\cdot)$ is the indicator function. At this point, we have the model $\hat{\theta}_{t-1}$ used for inference on tasks 1 to $t{-}1$, which we take as the starting point for training. After training with Eq. 1, we obtain the optimal parameters $\theta_{t}^{*}$ on the dataset $\mathcal{D}_{t}$:
@@
A natural idea is to combine models trained on different tasks. Specifically, $\hat{\theta}_{t-1}$, which be used as the inference parameter for these tasks, performs well on tasks $1$ to $t-1$, while $\theta_{t}^{*}$, which is optimized for task $t$, serves as the inference parameter for task $t$. Therefore, we can merge the two models. Here we use a simple convex combination:
@@
which means that taking the weighted average of $\hat{\theta}_{t-1}$ and $\theta_{t}^{*}$ is equivalent to scaling $\Delta\theta_{t}^{*}$. This scaling strategy means that it will not harm the parameters $\hat{\theta}_{t-1}$ for tasks 1 to $t-1$, but only by adjusting the task vector for task $t$ to reduce forgetting of old tasks. In our experiments 4.3, we found that this scaling has a slight impact on the knowledge of the new task.
@@
Ours (P&\&&M) $\theta_{t}^{*}=\theta_{t-1}+\Delta\theta_{t}^{*}$ $\hat{\theta}_{t}=\hat{\theta}_{t-1}+\alpha_{t}\Delta\theta_{t}^{*}$
@@ 3.2 A Closed-form Solution for Optimal Merge Coefficient
and use only the diagonal entries of $\hat{\mathbf{F}}(\theta)$. After completing the training of task $t$, we first compute the diagonal empirical Fisher matrix $\hat{\mathbf{F}}_{t}(\theta_{t}^{*})\approx{\mathbf{H}}_{t}(\theta_{t}^{*})$, then obtain $\alpha_{t}^{*}$ using Eq. 8, and finally compute the merged model $\hat{\theta}_{t}=\hat{\theta}_{t-1}+\alpha_{t}^{*}\Delta\theta_{t}^{*}$ as the inference parameters for tasks $1$ to $t$. Note that for each previous task $i<t$, the corresponding Fisher matrix $\hat{\mathbf{F}}_{i}(\theta_{i}^{*})$ is computed and stored at the time of training task $i$.
@@ 3.3 Train with Perturbation
We observe that Eq. 3.3 is effectively equivalent to applying a random perturbation to the model parameters during training, where the perturbation direction aligns with the task vector of task $t$. Specifically, the loss is evaluated under one of three perturbations: $\epsilon\Delta\theta_{t}$, $-\epsilon\Delta\theta_{t}$, or 00. Specifically, at each training step, we sample one of the three versions of the loss:
~$\tilde{\mathcal{L}}_{t}(\theta_{t})=\begin{cases}\mathcal{L}_{t}^{\text{ce}}(\theta_{t})&\text{with probability }p_{0},\\ \mathcal{L}_{t}^{\text{ce}}(\theta_{t}+\epsilon\Delta\theta_{t}^{*})&\text{with probability }p_{+},\\ \mathcal{L}_{t}^{\text{ce}}(\theta_{t}-\epsilon\Delta\theta_{t}^{*})&\text{with probability }p_{-}.\end{cases}$ (15)
@@
This sampling strategy reduces the forward cost per batch from 3× to 1×, without introducing bias into the gradient estimation. While variance may increase slightly, this technique enables us to scale P&\&&M to large models and datasets efficiently.
@@
15:     Estimate the diagonal empirical Fisher Information Matrix $\hat{\mathbf{F}_{t}}(\theta_{t}^{*})$ on task $\mathcal{T}_{t}$ using Eq. 10
@@ 3.4 LoRA-P&M
$\theta_{t}=\hat{\theta}_{t-1}+\sum_{p\in\mathcal{I}}\mathbf{A}_{t}^{(p)}\mathbf{B}_{t}^{(p)},$ (17)
In our experiments, we apply LoRA modules to the key and value projections, so that only the Fisher Information Matrices of these parameters need to be stored. The experiments comparing the memory and time overhead of our method with LoRA are provided in the Appendix  3. The overall procedure of the algorithm is presented in Algorithm  1.
@@ 4.1 Experimental Setups
Full Fine-Tuning 64.92$\pm$ 0.87 75.57$\pm$ 0.50 60.57$\pm$ 1.06 72.31$\pm$ 1.09 49.95$\pm$ 1.31 65.32$\pm$ 0.84
L2P wang2022learning 73.04$\pm$ 0.71 76.94$\pm$ 0.41 71.26$\pm$ 0.44 76.13$\pm$ 0.46 68.97$\pm$ 0.51 74.16$\pm$ 0.32
DualPrompt wang2022dualprompt 69.99$\pm$ 0.57 72.24$\pm$ 0.41 68.22$\pm$ 0.20 73.81$\pm$ 0.39 65.23$\pm$ 0.45 71.30$\pm$ 0.16
CODA-Prompt smith2023coda 76.63$\pm$ 0.27 80.30$\pm$ 0.28 74.05$\pm$ 0.41 78.14$\pm$ 0.39 69.38$\pm$ 0.33 73.95$\pm$ 0.63
HiDe-Prompt wang2023hierarchical 74.77$\pm$ 0.25 78.15$\pm$ 0.24 74.65$\pm$ 0.14 78.46$\pm$ 0.18 73.59$\pm$ 0.19 77.93$\pm$ 0.19
InfLoRA liang2024inflora 76.95$\pm$ 0.23 81.81$\pm$ 0.14 74.75$\pm$ 0.64 80.67$\pm$ 0.55 69.89$\pm$ 0.56 76.68$\pm$ 0.57
SD-LoRA (2025slora, ) 79.15$\pm$ 0.20 83.01$\pm$ 0.42 77.34$\pm$ 0.35 82.04$\pm$ 0.24 75.26$\pm$ 0.37 80.22$\pm$ 0.72
LoRA 71.22$\pm$ 1.47 78.15$\pm$ 1.08 65.72$\pm$ 0.75 76.14$\pm$ 0.96 56.35$\pm$ 0.80 71.08$\pm$ 1.04
@@ Appendix A Proof and Theoretical Details
where $\hat{\theta}_{t-1}$ denotes the inference-time model parameters after learning on tasks $1$ through $t-1$, and $\theta_{i}^{*}$ represents the optimal parameters for task $i$. The matrix $\mathbf{H}_{i}(\theta_{i}^{*})$ denotes the Hessian of the loss function for task $i$ evaluated at its optimum $\theta_{i}^{*}$. Substituting $\hat{\theta}_{t}=\hat{\theta}_{t-1}+\alpha_{t}\Delta\theta_{t}^{*},$ into the objective, we obtain:
@@
$J(\alpha_{t})=\sum_{i=1}^{t}\frac{1}{2}\Big{[}(\hat{\theta}_{t-1}-\theta_{i}^{*})^{\top}\mathbf{H}_{i}(\theta_{i}^{*})(\hat{\theta}_{t-1}-\theta_{i}^{*})+2\alpha_{t}(\hat{\theta}_{t-1}-\theta_{i}^{*})^{\top}\mathbf{H}_{i}(\theta_{i}^{*})\Delta\theta_{t}^{*}+\alpha_{t}^{2}(\Delta\theta_{t}^{*})^{\top}\mathbf{H}_{i}(\theta_{i}^{*})\Delta\theta_{t}^{*}\Big{]}.$ (21)
@@ Appendix B More Experimental Results and Details
Details of Datasets. ImageNet-R contains $200$ ImageNet [66] classes rendered in various artistic styles. ImageNet-A comprises $200$ classes featuring natural adversarial examples that are typically misclassified by standard ImageNet-trained models. DomainNet covers $345$ object categories across six distinct visual domains. CIFAR100 is a classic image classification dataset with $60,000$ images equally distributed across $100$ classes. CUB200 is a fine-grained bird classification dataset containing $11,788$ images over $200$ categories.
More results on DomainNet, ImageNet-A, CIFAR100, and CUB200. As in shown in Tab.  5 and  6 Across all four benchmarks, LoRA-P&M consistently outperforms both standard LoRA and the recent SD-LoRA. On challenging datasets such as ImageNet-A and DomainNet, it achieves the highest accuracy and AAA, surpassing SD-LoRA by up to +0.61% and +6.38%, respectively. Notably, on CIFAR100 and CUB200, LoRA-P&M matches or exceeds SD-LoRA’s accuracy while maintaining lower variance. These results highlight the robustness and generalization advantage of post-training task vector merging over both naive LoRA and structured dynamic variants like SD-LoRA.
Table 5: Performance comparison on ImageNet-A and DomainNet.
Method ImageNet-A 10 tasks DomainNet 5 tasks
Acc $\uparrow$ AAA $\uparrow$ Acc $\uparrow$ AAA $\uparrow$
L2P [33] 42.94 $\pm$ 1.27 51.40 $\pm$ 1.95 70.26 $\pm$ 0.25 75.83 $\pm$ 0.98
DualPrompt [58] 45.49 $\pm$ 0.96 54.68 $\pm$ 1.24 68.26 $\pm$ 0.90 73.84 $\pm$ 0.45
CODA-Prompt [32] 45.36 $\pm$ 0.78 57.03 $\pm$ 0.94 70.58 $\pm$ 0.53 76.68 $\pm$ 0.44
HiDe-Prompt [59] 42.70 $\pm$ 0.60 56.32 $\pm$ 0.40 72.20 $\pm$ 0.08 77.01 $\pm$ 0.04
InfLoRA [23] 49.20 $\pm$ 1.12 60.92 $\pm$ 0.61 71.59 $\pm$ 0.23 78.29 $\pm$ 0.50
SDLoRA [21] 55.96 $\pm$ 0.73 64.95 $\pm$ 1.63 72.82 $\pm$ 0.37 78.89 $\pm$ 0.50
SD-LoRA-RR [21] 55.59 $\pm$ 1.08 64.59 $\pm$ 1.91 72.58 $\pm$ 0.40 78.79 $\pm$ 0.78
SD-LoRA-KD [21] 54.24 $\pm$ 1.12 63.89 $\pm$ 0.58 72.15 $\pm$ 0.50 78.44 $\pm$ 0.66
LoRA 44.41 $\pm$ 0.57 56.76 $\pm$ 3.54 64.82 $\pm$ 2.30 75.28 $\pm$ 1.56
LoRA-P&\&&M 56.57 $\pm$ 0.78 65.35 $\pm$ 1.81 76.91 $\pm$ 0.53 85.27 $\pm$ 0.07
Table 6: Performance comparison on CIFAR100 and CUB200.
Method CIFAR100 10 tasks CUB200 10 tasks
Acc $\uparrow$ AAA $\uparrow$ Acc $\uparrow$ AAA $\uparrow$
L2P [33] 83.18 $\pm$ 1.20 87.69 $\pm$ 1.05 65.18 $\pm$ 2.49 76.12 $\pm$ 1.27
DualPrompt [58] 81.48 $\pm$ 0.86 86.41 $\pm$ 0.66 68.00 $\pm$ 1.06 79.40 $\pm$ 0.88
CODA-Prompt [32] 86.31 $\pm$ 0.12 90.67 $\pm$ 0.22 71.92 $\pm$ 0.33 78.76 $\pm$ 0.65
InfLoRA [23] 86.75 $\pm$ 0.35 91.72 $\pm$ 0.15 70.82 $\pm$ 0.23 81.39 $\pm$ 0.14
SD-LoRA [21] 88.01 $\pm$ 0.31 92.54 $\pm$ 0.18 77.48 $\pm$ 0.20 85.59 $\pm$ 0.44
LoRA 72.58 $\pm$ 1.57 80.64 $\pm$ 2.31 64.82 $\pm$ 2.30 75.28 $\pm$ 1.56
LoRA-P&\&&M 88.45 $\pm$ 0.35 92.89 $\pm$ 1.13 78.29 $\pm$ 0.50 83.39 $\pm$ 0.61
Hyperparameter ablation. We conduct ablation studies on the hyperparameters $p_{0}$ (probability of zero perturbation) and $\epsilon$ (perturbation magnitude) on the ImageNet-A dataset. As shown in Figure 3, we observe that setting $p_{0}=0.33$ and $\epsilon=0.5$ yields consistently good performance, though not necessarily optimal. To reduce tuning overhead and maintain consistency, we use $p_{0}=0.33$ and $\epsilon=0.5$ as default values for all experiments. Interestingly, when $p_{0}=0.33$, $p_{0}=p_{+}=p_{-}$.

### v4

@@ Abstract
Continual Learning (CL) aims to enable models to continuously acquire new knowledge from a sequence of tasks with avoiding the forgetting of learned information. However, existing CL methods only rely on the parameters of the most recent task for inference, which makes them susceptible to catastrophic forgetting. Inspired by the recent success of model merging techniques, we propose Perturb-and-Merge (P&M), a novel continual learning framework that integrates model merging into the CL paradigm to mitigate forgetting. Specifically, after training on each task, P&M constructs a new model by forming a convex combination of the previous model and the newly trained task-specific model. Through theoretical analysis, We minimize the total loss increase across all tasks and derive a closed-form solution for the merging coefficient under mild assumptions. To further improve the performance of the merged model, we observe that the degradation introduced during merging can be alleviated by a regularization term composed of the task vector and the Hessian matrix of the loss function. Interestingly, we show that this term can be efficiently approximated using second-order symmetric finite differences, and a stochastic perturbation strategy along the task vector direction is accordingly devised which incurs no additional forward or backward passes while providing an effective approximation of the regularization term. Finally, we combine P&M with LoRA, a parameter-efficient fine-tuning method, to reduce memory overhead. Our proposed approach achieves state-of-the-art performance on several continual learning benchmark datasets. The code is available at https://github.com/qhmiao/P-M-for-Continual-Learning.
@@ 3.1 Infer after Merging
where $\hat{p}_{y}$ is the predicted probability for the correct class $y$ and $\mathbb{I}(\cdot)$ is the indicator function. At this point, we have the model $\hat{\theta}_{t-1}$ used for inference on tasks 1 to t1t{-}1, which we take as the starting point for training. After training with Eq. 1, we obtain the optimal parameters $\theta_{t}^{*}$ on the dataset $\mathcal{D}_{t}$:
@@
A natural idea is to combine models trained on different tasks. Specifically, $\hat{\theta}_{t-1}$, which be used as the inference parameter for these tasks, performs well on tasks $1$ to t1t-1, while $\theta_{t}^{*}$, which is optimized for task $t$, serves as the inference parameter for task $t$. Therefore, we can merge the two models. Here we use a simple convex combination:
@@
which means that taking the weighted average of $\hat{\theta}_{t-1}$ and $\theta_{t}^{*}$ is equivalent to scaling $\Delta\theta_{t}^{*}$. This scaling strategy means that it will not harm the parameters $\hat{\theta}_{t-1}$ for tasks 1 to t1t-1, but only by adjusting the task vector for task $t$ to reduce forgetting of old tasks. In our experiments 4.3, we found that this scaling has a slight impact on the knowledge of the new task.
@@
Ours (P&\&M) $\theta_{t}^{*}=\theta_{t-1}+\Delta\theta_{t}^{*}$ $\hat{\theta}_{t}=\hat{\theta}_{t-1}+\alpha_{t}\Delta\theta_{t}^{*}$
@@ 3.2 A Closed-form Solution for Optimal Merge Coefficient
and use only the diagonal entries of $\hat{\mathbf{F}}(\theta)$. After completing the training of task $t$, we first compute the diagonal empirical Fisher matrix $\hat{\mathbf{F}}_{t}(\theta_{t}^{*})\approx-{\mathbf{H}}_{t}(\theta_{t}^{*})$, then obtain $\alpha_{t}^{*}$ using Eq. 8, and finally compute the merged model $\hat{\theta}_{t}=\hat{\theta}_{t-1}+\alpha_{t}^{*}\Delta\theta_{t}^{*}$ as the inference parameters for tasks $1$ to $t$. Note that for each previous task $i<t$, the corresponding Fisher matrix $\hat{\mathbf{F}}_{i}(\theta_{i}^{*})$ is computed and stored at the time of training task $i$.
@@ 3.3 Train with Perturbation
We observe that Eq. 3.3 is effectively equivalent to applying a random perturbation to the model parameters during training, where the perturbation direction aligns with the task vector of task $t$. Specifically, the loss is evaluated under one of three perturbations: $\epsilon\Delta\theta_{t}$, $\epsilon\Delta\theta_{t}$, or 0. Specifically, at each training step, we sample one of the three versions of the loss:
$\tilde{\mathcal{L}}_{t}(\theta_{t})=\begin{cases}\mathcal{L}_{t}^{\text{ce}}(\theta_{t})&\text{with probability }p_{0},\\ \mathcal{L}_{t}^{\text{ce}}(\theta_{t}+\epsilon\Delta\theta_{t}^{*})&\text{with probability }p_{+},\\ \mathcal{L}_{t}^{\text{ce}}(\theta_{t}-\epsilon\Delta\theta_{t}^{*})&\text{with probability }p_{-}.\end{cases}$ (15)
@@
This sampling strategy reduces the forward cost per batch from 3× to 1×, without introducing bias into the gradient estimation. While variance may increase slightly, this technique enables us to scale P&\&M to large models and datasets efficiently.
@@
15:  Estimate the Fisher Information Matrix $\hat{\mathbf{F}_{t}}(\theta_{t}^{*})$ on task $\mathcal{T}_{t}$ using Eq. 10 and Alg.  2
@@ 3.4 LoRA-P&M
$\theta_{t}\;=\;\theta_{t-1}\;+\;\bigoplus_{p\in\mathcal{P}}\Big(A_{t}^{(p)}B_{t}^{(p)}\Big),$ (17)
$\oplus$ denotes location-wise module insertion, not element-wise addition.
In our experiments, we apply LoRA modules to the key and value projections, so that only the Fisher Information Matrices of these parameters need to be stored. The experiments comparing the memory and time overhead of our method with LoRA are provided in the Appendix  B.3. The overall procedure of the algorithm is presented in Algorithm  1.
@@ 4.1 Experimental Setups
[…]
@@ 4.2 Main Results
More results on DomainNet, ImageNet-A, CIFAR100, and CUB200. As in shown in Tab.  4 and  5 Across all four benchmarks, LoRA-P&M consistently outperforms both standard LoRA and the recent SD-LoRA. On challenging datasets such as ImageNet-A and DomainNet, it achieves the highest accuracy and AAA, surpassing SD-LoRA by up to +0.61% and +6.38%, respectively. Notably, on CIFAR100 and CUB200, LoRA-P&M matches or exceeds SD-LoRA’s accuracy while maintaining lower variance. These results highlight the robustness and generalization advantage of post-training task vector merging over both naive LoRA and structured dynamic variants like SD-LoRA.
Table 4: Performance comparison on ImageNet-A and DomainNet.
Method ImageNet-A 10 tasks DomainNet 5 tasks
Acc $\uparrow$ AAA $\uparrow$ Acc $\uparrow$ AAA $\uparrow$
L2P (wang2022learning, ) 42.94 $\pm$ 1.27 51.40 $\pm$ 1.95 70.26 $\pm$ 0.25 75.83 $\pm$ 0.98
DualPrompt (wang2022dualprompt, ) 45.49 $\pm$ 0.96 54.68 $\pm$ 1.24 68.26 $\pm$ 0.90 73.84 $\pm$ 0.45
CODA-Prompt (smith2023coda, ) 45.36 $\pm$ 0.78 57.03 $\pm$ 0.94 70.58 $\pm$ 0.53 76.68 $\pm$ 0.44
HiDe-Prompt (wang2024hierarchical, ) 42.70 $\pm$ 0.60 56.32 $\pm$ 0.40 72.20 $\pm$ 0.08 77.01 $\pm$ 0.04
InfLoRA (liang2024inflora, ) 49.20 $\pm$ 1.12 60.92 $\pm$ 0.61 71.59 $\pm$ 0.23 78.29 $\pm$ 0.50
SDLoRA (2025slora, ) 55.96 $\pm$ 0.73 64.95 $\pm$ 1.63 72.82 $\pm$ 0.37 78.89 $\pm$ 0.50
SD-LoRA-RR (2025slora, ) 55.59 $\pm$ 1.08 64.59 $\pm$ 1.91 72.58 $\pm$ 0.40 78.79 $\pm$ 0.78
SD-LoRA-KD (2025slora, ) 54.24 $\pm$ 1.12 63.89 $\pm$ 0.58 72.15 $\pm$ 0.50 78.44 $\pm$ 0.66
LoRA 44.41 $\pm$ 0.57 56.76 $\pm$ 3.54 64.82 $\pm$ 2.30 75.28 $\pm$ 1.56
LoRA-P&\&M 56.57 $\pm$ 0.78 65.35 $\pm$ 1.81 76.91 $\pm$ 0.53 85.27 $\pm$ 0.07
Table 5: Performance comparison on CIFAR100 and CUB200.
Method CIFAR100 10 tasks CUB200 10 tasks
Acc $\uparrow$ AAA $\uparrow$ Acc $\uparrow$ AAA $\uparrow$
L2P (wang2022learning, ) 83.18 $\pm$ 1.20 87.69 $\pm$ 1.05 65.18 $\pm$ 2.49 76.12 $\pm$ 1.27
DualPrompt (wang2022dualprompt, ) 81.48 $\pm$ 0.86 86.41 $\pm$ 0.66 68.00 $\pm$ 1.06 79.40 $\pm$ 0.88
CODA-Prompt (smith2023coda, ) 86.31 $\pm$ 0.12 90.67 $\pm$ 0.22 71.92 $\pm$ 0.33 78.76 $\pm$ 0.65
InfLoRA (liang2024inflora, ) 86.75 $\pm$ 0.35 91.72 $\pm$ 0.15 70.82 $\pm$ 0.23 81.39 $\pm$ 0.14
SD-LoRA (2025slora, ) 88.01 $\pm$ 0.31 92.54 $\pm$ 0.18 77.48 $\pm$ 0.20 85.59 $\pm$ 0.44
LoRA 72.58 $\pm$ 1.57 80.64 $\pm$ 2.31 64.82 $\pm$ 2.30 75.28 $\pm$ 1.56
LoRA-P&\&M 88.45 $\pm$ 0.35 92.89 $\pm$ 1.13 78.29 $\pm$ 0.50 83.39 $\pm$ 0.61
Table 6: Acc comparison between global and per-module $\alpha$ strategies on ImageNet-R.
Method 5 tasks 10 tasks 20 tasks
Per-module $\alpha$ 80.53 76.82 74.68
Global $\alpha$ (ours) 80.88 78.48 74.13
Ablation on the Merging Coefficient Strategy. A potential concern is that using a uniform merging coefficient $\alpha$ across all parameters might be overly restrictive. To evaluate this, we compared our global-$\alpha$ strategy with a more fine-grained variant that learns separate $\alpha$ values for each LoRA module. The results on ImageNet-R are summarized in Table 6. The performance differences are negligible. We further observed that the learned per-module $\alpha$ values were highly similar to each other, suggesting that a single global coefficient is sufficient to capture the merging dynamics. Therefore, we adopt the global-$\alpha$ strategy for its simplicity, efficiency, and theoretical tractability. This finding aligns with prior research ilharco2023editing , where a unified coefficient was shown to preserve the structural integrity of the task vector $\Delta\theta_{t}$, which represents a coherent direction in parameter space. In contrast, a parameter-wise $\alpha$ may introduce inconsistent scaling and distort the internal correlations of $\Delta\theta_{t}^{\ast}$, thereby complicating the derivation of closed-form solutions.
@@ 5 Conclusion
Acknowledgment
Miao Zhang was partially sponsored by the National Natural Science Foundation of China under Grant 62306084 and U23B2051, Shenzhen College Stability Support Plan under Grant GXWD20231128102243003, and Shenzhen Science and Technology Program under Grant ZDSYS20230626091203008 and KJZD20230923115113026. Ziyue Qiao, School of Computing and Information Technology, Great Bay University, was supported by the National Natural Science Foundation of China (Grant No. 62406056) and the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515140114).
@@ Appendix A Proof and Theoretical Details
where $\hat{\theta}_{t-1}$ denotes the inference-time model parameters after learning on tasks $1$ through t1t-1, and $\theta_{i}^{*}$ represents the optimal parameters for task $i$. The matrix $\mathbf{H}_{i}(\theta_{i}^{*})$ denotes the Hessian of the loss function for task $i$ evaluated at its optimum $\theta_{i}^{*}$. Substituting $\hat{\theta}_{t}=\hat{\theta}_{t-1}+\alpha_{t}\Delta\theta_{t}^{*},$ into the objective, we obtain:
@@
$J(\alpha_{t})$ $=\sum_{i=1}^{t}\frac{1}{2}\Big[(\hat{\theta}_{t-1}-\theta_{i}^{*})^{\top}\mathbf{H}_{i}(\theta_{i}^{*})(\hat{\theta}_{t-1}-\theta_{i}^{*})$ (21)
$+2\alpha_{t}(\hat{\theta}_{t-1}-\theta_{i}^{*})^{\top}\mathbf{H}_{i}(\theta_{i}^{*})\Delta\theta_{t}^{*}+\alpha_{t}^{2}(\Delta\theta_{t}^{*})^{\top}\mathbf{H}_{i}(\theta_{i}^{*})\Delta\theta_{t}^{*}\Big].$ (22)
@@ Appendix B More Experimental Results and Details
B.1 Details of Datasets
ImageNet-R contains $200$ ImageNet [66] classes rendered in various artistic styles. ImageNet-A comprises $200$ classes featuring natural adversarial examples that are typically misclassified by standard ImageNet-trained models. DomainNet covers $345$ object categories across six distinct visual domains. CIFAR100 is a classic image classification dataset with $60,000$ images equally distributed across $100$ classes. CUB200 is a fine-grained bird classification dataset containing $11,788$ images over $200$ categories.
B.2 Hyperparameter ablation
We conduct ablation studies on the hyperparameters $p_{0}$ (probability of zero perturbation) and $\epsilon$ (perturbation magnitude) on the ImageNet-A dataset. As shown in Figure 3, we observe that setting $p_{0}=0.33$ and $\epsilon=0.5$ yields consistently good performance, though not necessarily optimal. To reduce tuning overhead and maintain consistency, we use $p_{0}=0.33$ and $\epsilon=0.5$ as default values for all experiments. Interestingly, when $p_{0}=0.33$, $p_{0}=p_{+}=p_{-}$.